# A parametric study of the pipeline hammer phenomenon in plastic Bingham slurry flows using the finite element method

### v1

@@ Abstract
We do a numerical study of the transient phenomenon in pipelines transporting plastic Bingham slurry flows, using a lowest-order finite element method (FEM). While most pipeline hammer studies focus on Newtonian fluids, the transient dynamics in Bingham fluids remains elusive and poorly afforded, even though this case has significant impact in industry like mining. A detailed parametric study asses the effects of the slurry yield stress and the valve closure times on both, pressure and velocity distributions along the pipeline, using an adaptive friction model to account for turbulent slurries. Results reveal that yield stress enhances flow resistance and accelerates pressure peak attenuation, underscoring the damping role of Bingham rheology compared to Newtonian flows. These insights emphasize the need for advanced FEM-based schemes in non-Newtonian shockwave modeling, with implications for industrial pipeline design and operational safety.
@@ 1 Introduction
Water-hammer is a common and critical phenomenon in many hydro-transport systems around the world. This problem is often caused by sudden operations in transport systems, such as a sudden closing or opening of flow control valves, pump or turbines power outages, or the pipeline ruptures, to name a few examples [1]. These events can generate intense pressure shockwaves traveling along the pipeline at very high speeds, covering long distances in quite short times. This feature gives very short time for a mechanical response to the pipeline walls, resulting in a significant increase in flow’s pressure and, in extreme cases, approaching to the limit mechanical strength of the pipeline. Then, this is a significant challenge for the operation of pipelines including system’s failure, equipment damage and operational inefficiencies, among other [2].
@@
Recent studies have addressed this phenomenon affording specific topics such as shear-thinning, shear-thickening or even tixotropy effects on shockwave propagation [14], missing an overall analysis about the influence of physical-mechanical parameters on the transient evolution on these flows. The complexity of the modeling of solid-liquid flows is usually related to the rheological character of the flow, often exhibiting a mechanical behavior similar to the plastic Bingham model. This model is particularly present for volume concentrations up to $C_{V}\approx 30\%$ [15]. A dramatic increase in this concentration could lead to hyperconcentrated regimes governed by nonlinear rheological models [16, 17, 18].
@@ 1.1 One dimensional models
One-dimensional (1D) models are among the most popular schemes for water-hammer analysis. Often based on the method of characteristics (MOC) (e.g. [1, 24]), they can efficiently simulate shockwave propagation in pipes accounting for wall shear stress and friction, simplified flow-structure interaction models, presence of bifurcations or valve closure models, among others. Recent progress have improved the accuracy and computational cost of these models allowing for the simulation of large hydraulic systems. This approach becomes interesting when dealing with mining suspensions, considering the difficulty of solving the coupled motion equations associated to the solid and liquid phases [27, 28], but also because of the dynamical relations involved on the interaction between both phases require previous empirical validation which is not always available. The continuous modeling of pipeline hammer incorporating a rheological model emerges then as a simple and low-cost alternative for predicting pressure impacts in a pipe.
@@
In our case, a methodological difference concerning the spatial discretization of the governing dynamics of the system was considered. The gold standard has relied on the method of characteristics [2, 24] with a structured 2D grid for space and time. Taking into account the consolidated computational power of the last decades, we consider instead a finite element approach which allows one not only to solve the classical conservation equations, but also to extend them to several other scenarios. With classical MOC, the hypothesis of quasi-linearity must be ensured, meaning that the equations terms must be either linear or depend solely on the unknown but not its derivatives.
@@ 2 Governing Equations
In this article, the pressure and velocity wave propagation after the closure of a control valve at the pipeline downstream section, is modeled as depicted in Figure 2. In this setup, an upstream head reservoir of height $H$ is connected to the pipe of inner diameter $D$ and total length $L$. A control valve is located at the downstream section of the system, which can close instantaneously ($T=0$) or gradually ($T>0$) depending on the operating conditions, where $T$ is the closure time. The condition $T=0$ is usually considered the most adverse scenario for pipeline safety. In this context, the one-dimensional (1D) averaged momentum and mass balance equations reads:
@@
$\frac{\partial p}{\partial t}+\rho_{m}c^{2}\frac{\partial u}{\partial x}&=0&\text{ in }(x,t)\in[0,L]\times[0,T],\\ \rho_{m}\left(\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}\right)+\frac{\partial p}{\partial x}+\mathcal{S}(u)&=0&\text{ in }(x,t)\in[0,L]\times[0,T],\\ u&=g(t)&\text{ on }x=L,~{}t\geq 0,\\ p&=P_{o}:=\rho_{m}gH&\text{ on }x=0,~{}t\geq 0,\\ p&=P_{o}\left(1-\frac{x}{L}\right)&\text{ at }t=0,~{}x\in[0,L],\\ u&=\frac{16}{f\widetilde{\text{Re}}}U_{o}&\text{ at }t=0,~{}x\in[0,L],\end{aligned}\right.$ (2.1)
@@
where $\eta$ is the fluid Bingham viscosity and $\rho_{m}$ stands for the fluid density, which is understood as a mixture density for a bi-phasic media, as it will be explained later on. The parameters $P_{o},U_{o}$ stand for a characteristic pressure and a characteristic velocity, respectively. Namely, we set $P_{o}=\rho_{m}gH$ and $U_{o}$ as the stationary laminar flow mean velocity before the valve closure:
$U_{o}:=\frac{P_{o}D^{2}}{32L\eta},$
The reader may notice in the fore-coming sections that some simulations will surpasses the laminar regime. We will still admit $U_{o}$ as a rough approximation for the 1D velocity initial condition. In addition, notice the corrector term sometimes called conductance, $16/(f\widetilde{\text{Re}})$, which accounts for the non-Newtonian behavior of the fluid. Notice as well that $P_{o}$ is chosen as the boundary condition for the pressure at the head tank (at $x=0$), whereas a linear stationary profile is used for the pipeline pressure distribution accordingly with the aforementioned initial condition for the velocity. Last, the valve closure model is described by a time-dependent function $g(t)$ at $x=L$.
The parameter $c$ represents the shockwave celerity. In addition, $\widetilde{\text{Re}}=\frac{\rho U_{o}D}{\eta}$ stands for a reference Reynolds number computed with the characteristic velocity. These scales will be used later on to formulate the dimensionless form of the governing equations. In addition, an isothermic assumption is considered for the fluid compressibility leading to a constant value for the pressure wave speed $c$. Therefore, the equations describe a fluid flow with mixture density $\rho_{m}$, axial velocity $u$, pressure $p$ and a reaction term denoted by $\mathcal{S}(u)$. This last term models the wall shear stress effect on the overall motion. Using the Fanning friction factor $f$ [32], the cross-section velocity $u$ and pipe diameter $D$ we can write $\mathcal{S}(u)$ as follows ([2], chapter 2):
$\mathcal{S}(u)=\frac{2f\rho_{m}|u|u}{D}$ (2.2)
The fluid rheology is implicit in this system through the friction coefficient. For instance, for Bingham flows the coefficient $f\text{Re}$ can be readily estimated from the following relations [33, 34]:
$f\text{Re}=\left\{\begin{aligned} 16+\frac{10.67+0.1414\left(\frac{\text{He}}{\text{Re}}\right)^{1.143}}{1+0.0149\left(\frac{\text{He}}{\text{Re}}\right)^{1.16}}\left(\frac{\text{He}}{4\text{Re}}\right),\quad&\text{Re}<Re_{c},\\ \text{Re}\left(\frac{10^{a}}{4\text{Re}^{0.193}}\right),\quad&\text{Re}\geq Re_{c}\end{aligned}\right.$ (2.3)
where $a=-1.47(1+0.146\exp(-2.9\cdot 10^{-5}\text{He}))$ is an empirical coefficient [15]. Notice that these relations recover the Newtonian behavior when $\text{He}=0$. The Reynolds (Re) and Hedstrom (He) numbers are introduced as follows [15]:
Re $=\frac{\rho Du}{\eta}$ (2.4)
He $=\frac{\rho D^{2}\tau_{y}}{\eta^{2}}$
@@
where $\tau_{y}$ is the yield stress and $\eta$ the Bingham viscosity. The reader should notice that the system initial condition will heavily depend on both, the Reynolds and the Hedstrom number. For turbulent regimes, higher friction factors will be computed, thus leading to $16/(f\text{Re})\ll 1$. A similar phenomena will be depicted for large Hedstrom numbers, where high yield-stresses will naturally lead a larger friction with the walls. In Figure 1, we show the starting conditions for friction and velocity of the system, for a range of Hedstrom numbers, using the parameter setup from 1. The reader should remark that the friction factor is a non-constant parameter in the simulation, thus increasing the overall numerical scheme as it is necessary to recompute the factor $f\text{Re}$ at each time-step, in each mesh node.
@@
Figure 1: Initial conditions for the slurry system according to the parameter setup of a copper mixture, detailed in table 1
2.1 Wave propagation on homogeneous mixtures
@@
where $\rho_{s},\rho_{f}$ denotes the solid and fluid densities, respectively and $C_{V}$ the solid volumetric concentration. Theoretical estimations of the pressure surge $\Delta P$ caused by a sudden change in the velocity regime in a pipeline it’s a longstanding engineering challenge. Joukowski was the first to determine the pressure rise generated in a rigid-walled pipe transporting a Newtonian liquid of density $\rho$ [35, 36], that is:
$\Delta{P}=\rho_{m}cU_{o}$ (2.6)
@@
$c=\sqrt{\frac{\frac{K_{f}}{\rho_{f}}}{1+\zeta}}$ (2.7)
where $K_{f}$ is the liquid bulk modulus. According to Young [37], for rigid walls pipelines we have $\zeta=0$, and Korteweg shown [38, 39] that $\zeta=1+\frac{K_{f}}{E}\frac{D}{e}$ for elastic walls tubes, where $E$ is the pipe elastic modulus, $D$ the inner diameter of the tube and $e$ the wall thickness. Other expressions considering pipe constraints can be found in Chaudhry [2]. The previous formulas can be adapted for estimating the wave speed celerity $c_{m}$ for homogeneous and heterogeneous solid-liquid mixtures. In the former, the expression reported by Young can be simply adapted as follows:
$c=\sqrt{\frac{K_{m}}{\rho_{m}}}$ (2.8)
@@
$c=\sqrt{\frac{\frac{K_{m}}{\rho_{m}}}{1+\frac{K_{m}}{E}\frac{D}{e}}}$ (2.9)
where $e$ stands for the wall thickness. Finally, another expression also valid for homogeneous mixtures and that considers the volumetric solid concentration $C_{V}$, reads [40, 41]:
$c=\sqrt{\frac{\frac{K_{f}}{\rho_{m}}}{1-C_{V}+\frac{K_{f}}{K_{s}}C_{V}+\frac{D}{e}\frac{K_{f}}{E}}}$ (2.10)
@@ 2.2 The transition velocity for Bingham flows
$\text{Re}_{c}=\frac{\text{He}}{X_{c}}\left(1-\frac{4}{3}X_{c}+\frac{1}{3}X_{c}^{4}\right)$ (2.11)
where $X_{c}=\tau_{y}/(\tau_{w})_{c}$, and $(\tau_{w})_{c}$ is the critical wall shear stress. Just at the laminar-transition point Hanks proposed:
@@
This means we can compute the critical velocity and therefore the transition number $\text{Re}_{c}$ fixing a Hedstrom number, thus allowing us to study several rheological behaviors of the flow. Remark as well that $X_{c}$ can be physically interpreted as a Bingham number $\text{Bi}=\tau_{y}/(\tau_{w})_{c}$ for a fully developed Bingham flow, which is the only regime in which this relation is valid. For the Newtonian case, the threshold value $\text{Re}_{c}=2100$ will be employed following [42].

### v2

@@ Abstract
We conduct a numerical study of the transient phenomenon in pipelines transporting plastic Bingham slurry flows, using a lowest-order finite element method (FEM). While most pipeline hammer studies focus on Newtonian fluids, the transient dynamics in Bingham fluids remains elusive and poorly afforded, despite their significant industrial impact, particularly in mining. A detailed parametric study assesses the effects of the slurry yield stress and the valve closure times on both pressure and velocity distributions along the pipeline, using an adaptive friction model to account for turbulent slurries. Results reveal that yield stress enhances flow resistance and accelerates pressure peak attenuation, underscoring the damping role of Bingham rheology compared to Newtonian flows. These insights emphasize the need for advanced FEM-based schemes in non-Newtonian shockwave modeling, with implications for industrial pipeline design and operational safety.
@@ 1 Introduction
Water-hammer is a common and critical phenomenon in many hydro-transport systems around the world. This problem is often caused by sudden operations in transport systems, such as a sudden closing or opening of flow control valves, pump or turbines power outages, or the pipeline ruptures, to name a few examples [1, 2, 3]. These events can generate intense pressure shockwaves traveling along the pipeline at very high speeds, covering long distances in quite short times. This feature gives very short time for a mechanical response to the pipeline walls, resulting in a significant increase in flow’s pressure and, in extreme cases, approaching to the limiting mechanical strength of the pipeline. Then, this is a significant challenge for the operation of pipelines including system failure, equipment damage and operational inefficiencies, among other [4].
@@
Recent studies have addressed this phenomenon affording specific topics such as shear-thinning, shear-thickening or even thixotropy effects on shockwave propagation [16], missing an overall analysis about the influence of physical-mechanical parameters on the transient evolution on these flows. The complexity of the modeling of solid-liquid flows is usually related to the rheological character of the flow, often exhibiting a mechanical behavior similar to the plastic Bingham model. This model is particularly present for volume concentrations up to $C_{V}\approx 30\%$ [17]. A dramatic increase in this concentration could lead to hyperconcentrated regimes governed by nonlinear rheological models [18, 19, 20].
@@ 1.1 One dimensional models
One-dimensional (1D) models are among the most popular schemes for water-hammer analysis. Often based on the method of characteristics (MOC) (e.g. [3, 26]), they can efficiently simulate shockwave propagation in pipes accounting for wall shear stress and friction, simplified flow-structure interaction models, presence of bifurcations or valve closure models, among others. Recent progress have improved the accuracy and computational cost of these models allowing for the simulation of large hydraulic systems. This approach becomes interesting when dealing with mining suspensions, considering the difficulty of solving the coupled motion equations associated to the solid and liquid phases [29, 30], but also because of the dynamical relations involved on the interaction between both phases require previous empirical validation which is not always available. The continuum modeling of pipeline hammer incorporating a rheological model emerges then as a simple and low-cost alternative for predicting pressure impacts in a pipe.
@@
In our case, a methodological difference concerning the spatial discretization of the governing dynamics of the system was considered. The gold standard has relied on the method of characteristics [4, 26] with a structured 2D grid for space and time. Taking into account the consolidated computational power of the last decades, we consider instead a finite element approach which allows one not only to solve the classical conservation equations, but also to extend them to several other scenarios.
While the classical Method of Characteristics (MOC) has long served as the gold standard, particularly in conjunction with structured 2D grids for space and time [4, 26], its applicability is inherently limited to first-order hyperbolic PDEs, where nonlinearities appear only in the unknowns and not in their derivatives. This limitation becomes significant when addressing higher-order or fully nonlinear dynamics, such as those arising from more advanced viscous or constitutive models, where the characteristic structure may break down and numerical artifacts can appear without convergence guarantees. In contrast, the finite element method (FEM) provides a more general and robust framework that naturally accommodates such complex formulations. By allowing flexible discretization in both space and time and supporting weak formulations of arbitrary order, FEM enables the extension of our modeling approach beyond classical conservation laws to a broader class of non-Newtonian and nonlinear flow regimes. For example, extending the shear model using the admittedly simple power-law approach [34] results in a second-order nonlinear PDE, for which characteristic curves cannot be consistently defined. Similarly, elasto-viscoplastic behavior can be analyzed with nonlinear models such as those explored in [35], where additional transport equations are incorporated into the overall dynamics.
@@ 2 Governing Equations
[…]
@@
$\frac{\partial p}{\partial t}+\rho_{m}c^{2}\frac{\partial u}{\partial x}&=0&\text{ in }(x,t)\in[0,L]\times[0,T],\\ \rho_{m}\left(\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}\right)+\frac{\partial p}{\partial x}+\mathcal{S}(u)&=0&\text{ in }(x,t)\in[0,L]\times[0,T],\\ u&=\frac{16}{f(\widetilde{\text{Re}},\text{He})\widetilde{\text{Re}}}g(t)&\text{ on }x=L,~t\geq 0,\\ p&=P_{o}:=\rho_{m}gH&\text{ on }x=0,~t\geq 0,\\ p&=P_{o}\left(1-\frac{x}{L}\right)&\text{ at }t=0,~x\in[0,L],\\ u&=\frac{16}{f(\widetilde{\text{Re}},\text{He})\widetilde{\text{Re}}}U_{o}&\text{ at }t=0,~x\in[0,L],\end{aligned}\right.$ (2.1)
@@
where $f(\text{Re},\text{He})$ is a non-constant friction factor, i.e., it varies in space and time with the local velocity, and $\rho_{m}$ stands for the mixture density, as it will be explained later in Section 2.1. The parameters $P_{o},U_{o}$ stand for characteristics pressure and velocity scales, respectively. Namely, we set $P_{o}=\rho_{m}gH$ and $U_{o}$ as the stationary laminar flow mean velocity before the valve closure:
$U_{o}:=\frac{P_{o}D^{2}}{32L\eta},$ (2.2)
where $\eta$ is the fluid Bingham viscosity. In other words, we set $U_{o}$ as the Hagen-Poiseuille formula for average flow velocity in a circular pipe.
In addition, $\widetilde{\text{Re}}=\frac{\rho U_{o}D}{\eta}$ stands for a reference (and constant) Reynolds number computed with the characteristic velocity. On the other hand, the instantaneous Reynolds (Re) and Hedstrom (He) numbers are introduced as follows [17]:
Re $=\frac{\rho Du}{\eta},$ (2.3)
He $=\frac{\rho_{m}D^{2}\tau_{y}}{\eta^{2}},$
@@
where $\tau_{y}$ is the fluid yield stress. It’s worth noting that the system initial condition will heavily depend on both, the Reynolds and the Hedstrom number. For turbulent regimes, higher friction factors will be computed, thus leading to $16/(f\text{Re})\ll 1$. A similar phenomena will be depicted for large Hedstrom numbers, where high yield-stresses will naturally lead a larger friction with the walls. In Figure 2, we show the starting conditions for friction and velocity of the system, for a range of Hedstrom numbers, using the parameter setup from Tables 1, 2 and 3, for a 30% concentration copper slurry. The reader should note that the friction factor is a non-constant parameter in the simulation, thus increasing the overall numerical complexity as it is necessary to recompute the factor $f\text{Re}$ at each time-step, in each mesh node.
The reader will notice in the following sections that some simulations extend beyond the laminar regime. Nevertheless, we continue to use $U_{0}$ as a rough approximation for the initial 1D velocity field. To account for the non-Newtonian behavior of the fluid, we include a corrector term—sometimes referred to as conductance [32]—given by $16/(f\widetilde{\text{Re}})$. Regarding the pressure field, $P_{0}$ is imposed as a boundary condition at the head tank, located at $x=0$ (see Figure 1), while the pressure distribution along the pipeline follows a linear steady-state profile consistent with a fully developed Poiseuille pipe flow, given by $p(x)=P_{0}(1-x/L)$. Finally, the valve closure at the downstream end ($x=L$) is modeled through a time-dependent function $g(t)$.
The parameter $c$ represents the shockwave celerity. These scales will be used later on to formulate the dimensionless form of the governing equations. In addition, an isothermic assumption is considered for the fluid compressibility leading to a constant value for the pressure wave speed $c$. Therefore, the equations describe a fluid flow with mixture density $\rho_{m}$, axial velocity $u$, pressure $p$ and a reaction term denoted by $\mathcal{S}(u)$. This last term models the wall shear stress effect on the overall motion. Using the Fanning friction factor $f$ [36], the cross-section velocity $u$ and pipe diameter $D$ we can write $\mathcal{S}(u)$ as follows ([4], chapter 2):
$\mathcal{S}(u)=\frac{2f(\text{Re},\text{He})\rho_{m}|u|u}{D},$ (2.4)
which is a quasi-steady approach. This represents a simplified formulation of the wall shear stress model [6, 37, 38], as unsteady velocity profiles have been shown to differ significantly from those in the steady-state regime [39], regardless of the fluid rheology [16]. For example, in [40], a comprehensive set of analytic expressions for unsteady friction losses is studied in the Newtonian scenario. Similarly, in [41] a comparative study involving six different models assesses them not only based on how they reproduce the wave propagation, but also under an entropy-based metric. In [42], a convolution-based model is used for the friction losses, as is usually done with damping models where the equations are more easily handled in the frequency domain.
Furthermore, recent studies [43] have integrated unsteady friction losses in contemporary optimization problems where the governing dynamics are introduced in a cost function where the state is approximated with a neural network [44]. This approach with physics-informed neural networks (PINNs) relates to a class of inverse problems where the physical information of the phenomena can be entangled with sensor data [45, 46], which is another future branch to be studied by the authors.
The quasi-steady approach is broadly used by the community and is also present in other computer codes [47, 48], so it is still an open subject to be tackled in future works. We also remark that $\mathcal{S}(u)$ will be computed locally in time and space, meaning that the friction factor will vary as a function of the fluid Reynolds number and the fluid Hedstrom number.
The fluid rheology is implicit in this system through the Fanning friction coefficient. For instance, for Bingham flows the coefficient $f\text{Re}$ can be readily estimated from the following relations [49, 50, 51]:
$f(\text{Re},\text{He})\text{Re}=\left\{\begin{aligned} 16+\frac{10.67+0.1414\left(\frac{\text{He}}{\text{Re}}\right)^{1.143}}{1+0.0149\left(\frac{\text{He}}{\text{Re}}\right)^{1.16}}\left(\frac{\text{He}}{4\text{Re}}\right),\quad&\text{Re}<\text{Re}_{c},\\ \text{Re}\left(\frac{10^{a}}{4\text{Re}^{0.193}}\right),\quad&\text{Re}\geq\text{Re}_{c},\end{aligned}\right.$ (2.5)
[…]
@@
Figure 2: Initial conditions for the slurry system according to the parameter setup of a copper mixture, detailed in Tables 1, 2 and 3
2.1 Shockwave propagation in homogeneous mixtures
@@
where $\rho_{s},\rho_{f}$ denotes the solid and fluid densities, respectively and $C_{V}$ the solid volumetric concentration. Theoretical estimations of the pressure surge $\Delta P$ caused by a sudden change in the velocity regime in a pipeline is a longstanding engineering challenge. Joukowski was the first to determine the pressure rise generated in a rigid-walled pipe transporting a Newtonian liquid of density $\rho_{m}$ [52], that is:
$\Delta{P}=\rho_{m}cU_{o},$ (2.7)
@@
$c=\sqrt{\frac{\frac{K_{f}}{\rho_{f}}}{1+\zeta}},$ (2.8)
where $K_{f}$ is the liquid bulk modulus. According to Young [53], for rigid walls pipelines we have $\zeta=0$, and Korteweg showed [54, 55] that $\zeta=1+\frac{K_{f}}{E}\frac{D}{e}$ for elastic walls tubes, where $E$ is the pipe elastic modulus, $D$ the inner diameter of the tube and $e$ the wall thickness. Other expressions considering pipe constraints can be found in [4, 56]. The previous formulas can be adapted for estimating the wave speed celerity $c_{m}$ for homogeneous and heterogeneous solid-liquid mixtures. In the former, the expression reported by Young can be simply adapted as follows:
$c=\sqrt{\frac{K_{m}}{\rho_{m}}},$ (2.9)
@@
$c=\sqrt{\frac{\frac{K_{m}}{\rho_{m}}}{1+\frac{K_{m}}{E}\frac{D}{e}}}.$ (2.10)
Finally, another expression also valid for homogeneous mixtures read [57, 58]:
$c=\sqrt{\frac{\frac{K_{f}}{\rho_{m}}}{1-C_{V}+\frac{K_{f}}{K_{s}}C_{V}+\frac{D}{e}\frac{K_{f}}{E}}}.$ (2.11)
@@ 2.2 The transition velocity for Bingham flows
$\text{Re}_{c}=\frac{\text{He}}{X_{c}}\left(1-\frac{4}{3}X_{c}+\frac{1}{3}X_{c}^{4}\right),$ (2.12)
where $X_{c}=\tau_{y}/(\tau_{w})_{c}$ and $(\tau_{w})_{c}$. Just at the laminar-transition point Hanks and Pratt [59] demonstrated that Hedstrom number could be correlated with the critical Reynolds number, fixing the onset of turbulence. This correlation can be written as follows:
@@
This means we can compute the critical velocity and therefore the transition number $\text{Re}_{c}$ fixing a Hedstrom number, thus allowing us to study several rheological behaviors of the flow. Remark as well that $X_{c}$ can be physically interpreted as a Bingham number $\text{Bi}=\tau_{y}/(\tau_{w})_{c}$ for a fully developed Bingham flow, which is the only regime in which this relation is valid. Just for reference, for the Newtonian case, the threshold value is $\text{Re}_{c}=2,100$ [60], but in a slurry this threshold can be higher.